In triangle PQR, A is the point of intersection of all the altitudes and B is the point of intersection of all the angle bisectors of the triangle. If ∠PBR = 105°, then what is the value of ∠PAR (in degrees)?

Grade:12th pass

2 Answers

Arun
5 years ago
Dear student...this is very simple question. Just try to draw a figure, and then find angle. Angle PAR will be 60°.....hope it helps
Shashi Kamal
5 years ago
Since Point B is the intersection point of angle bisectors then it must be incentre. /_PBR=90°+Q/2 so Q=30°. Also A is the intersection point of altitudes so it will be Orthocentre then PAR=(180-30)°=150°